OpenStudy (anonymous):

Need help with integrals The assignment is to find the area of: y = 6x - 3x² +2 I get the antiderivative: ((6x²)/2) - ((3x³)/2) + 2x But, when putting in the value 5 for x, I get -40... so, not the correct area

OpenStudy (agent0smith):

You made a mistake... \[\large \int\limits\limits_{a}^{b} (6x - 3x² +2 )dx =\left[ \frac{ 6x^2 }{ 2 } - \frac{3x^3 }{ 3 } +2x \right]_{a}^{b}\]

OpenStudy (agent0smith):

Increase the power by one, and divide by the *new* power. \[\large \int\limits\limits_{}^{} x^a dx = \frac{ x^{a+1} }{ a+1 } +C\] or with limits \[\large \int\limits\limits_{c}^{d} x^a dx = \left[ \frac{ x^{a+1} }{ a+1 }\right]_{c}^{d}\]

OpenStudy (anonymous):

Ah yes... even so the answer is wrong though :/ ((6x²)/2) - ((3x³)/3) + 2x

OpenStudy (anonymous):

Solved! Thanks